Arc cosx + arc cosy + arc cosz = π and if; x+y+z=3/2 then show that x=y=z
1
Expert's answer
2018-02-09T05:14:07-0500
Range of arccos(a) is [0,π ], the domain consists of values a from[-1,1]. But π = π/ 3+π/3+π/3 and arccos(π/3) = 1/2,in this case if x+y+z = 3/2, then x=y=z.